Question text Use the following information to answer Questions 6 and 7: •An analyst gathered the following information regarding Alturius Inc: •Current market price per share = $29.48 •Current year EPS = $3.18 •Current year dividend per share = $1.272 •Required rate of return on equity = 12% Dividends are expected to grow at a rate of 6.5% forever.

Question: The company's justified leading P/E ratio is closest to: Select one: a. 7.75 b. 7.27 c. 10.91

Question text Based on the justified trailing P/E ratio, the stock is most likely: Select one: a. Undervalued. b. Overvalued. c. Fairly valued.
Answer #1

Price of Stock = Current Year Dividend * (1 + Growth rate) / (required rate - Growth Rate)

Price of Stock = 1.272 * 1.065 / (12% - 6.50%)

Price of Stock = 1.35468 / 5.50%
Price of Stock = $24.63

Justified leading P/E ratio = Price / Earnings = $24.63 / $3.18 = 7.75 times Option A

the stock is most likely Undervalued Option A